You purchase 120 shares for $70 a share ($8,400), and after a year the price rises to $80. Calculate the percentage return on your investment if you bought the stock on margin and the margin requirement was (ignore commissions, dividends, and interest expense):

  1. 15 percent. Round your answer to one decimal place.

      %

  2. 55 percent. Round your answer to one decimal place.

      %

  3. 80 percent. Round your answer to one decimal place.

      %
Answer #1

Total Value at purchase = 120 x 70 = 8,400

Total Value after a year = 120 x 80 = 9,600

a) 15% margin, equity = 8,400 x 15% = 1,260, loan = 8,400 - 1,260 = 7,140

% Return = (9,600 - 7,140) / 1,260 = 95.24%

b) 55% margin, equity = 8,400 x 55% = 4,620, loan = 3,780

% Return = (9,600 - 3,780) / 4,620 = 25.97%

c) equity = 80% x 8400 = 6,720, loan = 1,680

% Return = (9600 - 1680) / 6720 = 17.86%